# Voting by mail: a Markov chain model for managing the security risks of absentee voting

### v1

@@ 6.2.1 X9: Malicious attack to challenge the signature of a valid ballot
Malicious $X9$ attack captures bad actors in an election office erroneously rejecting marked ballots by challenging their signatures. This scenario, with a base arc transition probability of 0.055, is set to affect roughly one of 19 precincts in Milwaukee. Referencing Fig. 2, malicious attack $X9$ impacts process state $VII$, where ballots are collected and verified by the election office. On a specific day, this could lead to a significant number of ballots being improperly discarded.
@@ 6.2.2 X13: Malicious attack discard ballot in transit via USPS
Malicious $X{13}$ attack represents bad actors in a postal office who discard completed ballots at the sorting or storage point. Referring back to Fig. 2, we see that malicious attack $X{13}$ directly impacts process state $VI$ where the voter returns the completed ballot via USPS. In Milwaukee County, a large-scale attack affects one of the 36 USPS offices for a day [12]. Consequently, there is a probability of 0.033 that ballots in process state $VI$ are impacted by the $X{13}$ attack.
@@
6.2.3 Masquerade Attacks, Ballot Modifying
We then consider malicious attack $X{29}$, which alters voters’ original ballots, impacting process states $II$. The consequence of this attack leads to altered ballots that are both counted and uncounted (final ballot states (C,A) and (NC,A)). The arc probability strengths from attack $X{9}$ are reflected here, since there exists no known reference of an executed masquerade attack.
@@ 6.2.3 Masquerade Attacks, Ballot Modifying
Vary attack strength X29 with attacks in interval 2 40 0.025 326,097 560 45 175 20,522 50
40 0.050 326,007 560 90 175 20,517 99
40 0.075 325,918 560 134 174 20,512 149
40 0.100 325,828 560 179 174 20,507 198
@@ 6.3 Worst-Case Scenario Modeling
We build on the attack scenarios in Section 6.2 to create a “worst-case scenario” to evaluate various policy implementations in the VBM process to severe conditions as wel as the impact on mitigations. The worst-case scenario features the simultaneous execution of the following three malicious attacks within the election cycle: $X{13}$ on day 37, $X{29}$ on day 7, and $X{9}$ on day 38. The following arc probabilities are all set to 0.10 on these days: arc from node $IV$ to $X{13}$ on day 37; arc from node $II$ to $X{29}$ on day 7; arc from node $VII$ to $X{9}$. We also consider a sensitivity analysis that evaluates the impact of mitigation strength and associated strength against the worst-case attack scenario. Referencing Table 3 in Section 3.4, four mitigations counter the attacks in the worst-case scenario (see Table 3):
@@
[…]
  • $X{29}$ is countered by ($M3,M4$)
We first vary drop box availability ($M7$) from its base value of $0.520$ using an availability range from $0.1$ to $0.95$. We do not consider $M7$ to have zero availability, since voters can return their absentee ballots directly to their local municipality election administration office. The remaining mitigations ($M6,M4,M3$) vary in availability, ranging from $0.01$ to $0.95$ to capture a wide range of operating conditions based on current federal laws and reports of mitigation effectiveness [2].
Next, we vary mitigation strength. Drop box $M7$ availability significantly impacts the number of ballots in the final ballot states. Increasing the availability of $M7$ to $0.95$ increases the counted ballots (C,U) to 324,867, which is higher the 324,792 (C,U) ballots in the baseline scenario without malicious attacks. Reducing $M7$ availability to less than 0.10 results in 17,776 (NC,U) ballots, the highest across all worst-case scenarios. However, M7 does not counter masquerade attacks to reduce the number of altered ballots (C,A) or (NC,A).
The following eight rows in Table 11 examine the availability of mitigation $M6$ associated with in-person absentee. Changes in $M6$ availability has substantially less impact on final ballot states than mitigation $M7$ availability. Similar to $M7$, the availability of mitigation $M6$ fails to counter masquerade attacks by reducing the number of altered ballots (C,A) or (NC,A).
Next, we examine varying mitigations $M3$ and $M4$ simultaneously to reflect their interdependency. We find that deploying automatic ballot notifications ($M3$) with the same availability as $M4$, at $0.90$, results in 325,584 counted (C,U) ballots, an increase of 9,742 ballots as compared to the baseline, worst-case performance in the first row of Table 11. An $M3$ availability of 0.90 results in 62 expected (C,A) ballots. Lastly, since $M3$ and $M6$ can function interdependently, we examine the impact of simultaneously varying $M3$ and $M6$. When deployed at 0.90 availability, we observe 325,485 (C,U) ballots, similar to the previous observations. Additionally, we observe fewer ballots not returned, with 17,243 (NC,NR) ballots compared with 20,371 (NC,NR) ballots in the Milwaukee baseline without malicious attacks. We conclude that in-person absentee voting gives voters a powerful recourse option to counter malicious and non-malicious attacks.
@@
Next, we study the sensitivity of the DTMC model mitigation strength. The two final ballot states we study are the desired final ballot state (C,U) and the unreturned ballots (NC,NR) since these two final ballot states are the most impacted by mitigations. However, other ballot states are impacted by mitigations to a lesser degree. We perform a one-way sensitivity analysis by changing each mitigation strength by ±$\pm 0.01$ from its baseline value (see Table 2). Figure 5 shows a tornado graph illustration of the results. The blue half of Fig. 5 on the left represents the difference in the number of correctly counted and unaltered ballots when the mitigation strength decreases. The red half is the difference in the number of ballots that are not counted and not returned, which occurs when mitigation strength increases. The tornado graph shows that the DTMC model is most sensitive to mitigations $M7$ and $M3$. In contrast, mitigations $M4$ and $M6$ (independently) do not significantly affect the outputs in a significant manner. However, their impact can be larger when their availability is considered within the context of other mitigations (see Table 11).
@@ 7 Conclusions
Singular mitigations, such as drop boxes, can bolster the VBM system, while other mitigations must be combined to reduce risk. A crucial observation is that mitigation $M3$ (dependent on $M4$ or $M6$) is the only mitigation capable of effectively countering ballot-altering malicious attacks. The variation of each mitigation policy set changes the VBM process’s resilience to the worst-case scenario. These changes mark quantifiable metrics for evaluating VBM security and highlight the need for an integrated and adaptive mitigation approach to ensure the integrity of the VBM process.
@@
Additionally, ballot tracking programs (M3) can significantly affect VBM performance. Both drop boxes and ballot tracking programs provide broad protection against attacks. Two notable examples of ballot tracking programs that are currently available are BalloTrax and BallotScout. BalloTrax allows voters to track the status of their mail-in ballots, providing real-time updates when ballots are collected, received, and accepted by election offices [9]. Similarly, BallotScout integrates with USPS to provide voters with detailed information on their ballots’ whereabouts throughout the election process [7].
There are several limitations of our study that provide insight into topics for future research. First, this analysis considers malicious attacks that occur on a single-day and is limited by a lack of historical data regarding malicious attacks. Future research could diversify attack scenarios and incorporate a broader range of data to enhance the model’s sensitivity to resource constraints and varying electoral processes. Additionally, future research could account for resource constraints and system congestion.

### v2

@@ 6.2.1 X9: Malicious attack to challenge the signature of a valid ballot
Malicious $X9$ attack captures bad actors in an election office erroneously rejecting marked ballots by challenging their signatures. This scenario, with a base arc transition probability of 0.055, is set to affect roughly one of 19 precincts in Milwaukee. Referencing Figure 2, malicious attack $X9$ impacts process state $VII$, where ballots are collected and verified by the election office. On a specific day, this could lead to a significant number of ballots being improperly discarded.
@@ 6.2.2 X13: Malicious attack discard ballot in transit via USPS
Malicious $X{13}$ attack represents bad actors in a postal office who discard completed ballots at the sorting or storage point. Referring back to Figure 2, we see that malicious attack $X{13}$ directly impacts process state $VI$ where the voter returns the completed ballot via USPS. In Milwaukee County, a large-scale attack affects one of the 36 USPS offices for a day [13]. Consequently, there is a probability of 0.033 that ballots in process state $VI$ are impacted by the $X{13}$ attack.
@@
6.2.3 X29: Masquerade Attacks, Ballot Modifying
We then consider the malicious attack $X{29}$, which alters voters’ original ballots and directly impacts process states $II$. The consequence of this attack leads to altered ballots that are both counted and uncounted (final ballot states (C,A) and (NC,A)). The arc probability strengths from attack $X{9}$ are reflected here, since there exists no known reference of an executed masquerade attack.
@@ 6.2.3 X29: Masquerade Attacks, Ballot Modifying
Vary attack strength X29 with attacks in interval 2 40 0.025 324,626 559 45 173 20,444 50
40 0.050 324,536 559 90 173 20,439 90
40 0.075 324,447 559 134 172 20,435 149
40 0.100 324,358 559 179 172 20,430 198
@@
6.2.4 Attack timing
We examine the impact of attack timing on three malicious attack scenarios: $X9$, $X13$, and $X29$. To do so, we vary the day of each attack from Day 7 to Day 49 and evaluate the change in the expected number of returned ballots that are counted and unaltered (C,U) compared to the baseline (see Table 6.2).
Figure 5 illustrates the deviations in the expected number of (C,U) ballots associated with each attack scenario as a function of the day of the attack. All values are negative, indicating that each attack reduces the number of ballots in the preferred (C,U) state. Attack $X9$, which models the false rejection of valid ballots at the election office (process state VII), experiences its largest deviation in (C,U) ballots when the attack is initiated on day 38. This attack peak aligns with a surge of ballots routed to the election office via early voting (M6), which begins on day 36 and introduces a direct arc from process state II to state VII, in addition to those being returned by mail or drop box. Similarly, attack $X13$, representing the loss of ballots in USPS transit, peaks on day 37 in terms of its deviation in the expected number of (C,U) ballots. This deviation increases due to ballot return reminders (M5) that start to alert voters on Day 36. This results in an increase in the flow of ballots in transit, which results in more ballots that are available to be maliciously stolen.
Attack $X29$, a masquerade attack involving stolen or pre-marked ballots (state II), experiences its largest deviation in (C,U) ballots earlier in the election cycle on day 7, which results in 706 fewer (C,U) ballots. This occurs due to the initial surge in ballot issuance and early returns at this time. A secondary deviation of 614 (C,U) ballots occurs on day 36, coinciding with an increase in voters returning ballots due to automatic ballot reminders (M5). These results indicate that attack timing–—specifically its alignment with the flow of ballots–—is crucial for determining the consequences of an attack. The consequences of an attack corresponds to its interaction with process dynamics during periods of high ballot flow.
Figure 5: Deviation in counted and unaltered (C, U) ballots from the baseline under three moderate malicious attack scenarios ($X9$, $X13$, and $X29$).
@@ 6.3 Worst-Case Scenario Modeling
We build on the attack scenarios in Section 6.2.4 to create a “worst-case scenario” to evaluate various policy implementations in the VBM process under severe conditions as well as the impact on mitigations. The worst-case scenario involves the simultaneous execution of three malicious attacks within the 49-day election cycle. Figure 5 illustrates the highest impact for each attack as: $X{9}$ on day 38, $X{13}$ on day 37, and $X{29}$ on day 7. The following arc probabilities are all set to 0.10 on these days: arc from node $IV$ to $X{13}$ on day 37; arc from node $II$ to $X{29}$ on day 7; arc from node $VII$ to $X{9}$. We also consider a sensitivity analysis that evaluates the impact of mitigation strength and associated strength against the worst-case attack scenario. Referencing Table 3 in Section 3.4, four mitigations counter the attacks in the worst-case scenario (see Table 3):
  • $X{9}$ is countered by automatic ballot notifications M3, replacement ballots M4, and early voting M6,
  • $X{13}$ is countered by automatic ballot notifications M3, replacement ballots M4, early voting M6, and drop boxes M7,
  • $X{29}$ is countered by automatic ballot notifications M3 and replacement ballots M4.
We first vary drop box ($M7$) availability from its base value of 0.5200.5200.520 using an availability range from 0.10.10.1 to 0.950.950.95. We do not consider drop boxes to have zero availability, since voters can return their absentee ballots directly to their local municipality election administration office. The remaining mitigations ($M6,M4,M3$) vary in availability, ranging from 0.010.010.01 to 0.950.950.95 to capture a wide range of operating conditions based on current federal laws and reports of mitigation effectiveness [2].
Next, we vary mitigation strength. Drop box $M7$ availability significantly impacts the number of ballots in the final ballot states. Increasing the availability of $M7$ to 0.950.950.95 increases the counted ballots (C,U) to 324,867, which is higher the 324,792 (C,U) ballots in the baseline scenario without malicious attacks. Reducing $M7$ availability to less than 0.10 results in 17,776 (NC,U) ballots, the highest across all worst-case scenarios. However, M7 does not counter masquerade attacks to reduce the number of altered ballots (C,A) or (NC,A).
The following eight rows in Table 11 examine the availability of early voting $M6$. Changes in early voting $M6$ availability has substantially less impact on final ballot states than drop boxes $M7$ availability. Similar to drop boxes $M7$, the availability of early voting $M6$ fails to counter masquerade attacks by reducing the number of altered ballots (C,A) or (NC,A).
@@
Next, we study the sensitivity of the DTMC model mitigation strength. The two final ballot states we study are the desired final ballot state (C,U) and the unreturned ballots (NC,NR) since these two final ballot states are the most impacted by mitigations. However, other ballot states are impacted by mitigations to a lesser degree. We perform a one-way sensitivity analysis by changing each mitigation strength by ±0.01$\pm$ 0.01± 0.01 from its baseline value (see Table 2). Figure 6 shows a tornado graph illustration of the results. The blue half of Figure 6 on the left represents the difference in the number of correctly counted and unaltered ballots when the mitigation strength decreases. The red half is the difference in the number of ballots that are not counted and not returned, which occurs when mitigation strength increases. The tornado graph shows that the DTMC model is most sensitive to drop boxes $M7$ and automatic ballot notifications $M3$. In contrast, replacement ballots $M4$ and early voting $M6$ (independently) do not significantly affect the outputs in a significant manner. However, their impact can be larger when their availability is considered within the context of other mitigations (see Table 11). These results demonstrate that modeling attack timing in conjunction with mitigation availability provides crucial insight into the consequence of each attack.
@@ 7 Conclusions
Singular mitigations, such as drop boxes, can bolster the VBM system, while other mitigations must be combined to reduce risk. A crucial observation is that automatic ballot notifications (dependent on replacement ballots or early voting) is the only mitigation capable of effectively countering ballot-altering malicious attacks. The variation of each mitigation policy set changes the VBM process’s resilience to the worst-case scenario. These changes mark quantifiable metrics for evaluating VBM security and highlight the need for an integrated and adaptive mitigation approach to ensure the integrity of the VBM process.
@@
Additionally, automatic ballot tracking can significantly affect VBM performance. Both drop boxes and automatic ballot tracking provide broad protection against attacks. Two notable examples of ballot tracking programs that are currently available are BallotTrax and BallotScout. BallotTrax allows voters to track the status of their mail-in ballots, providing real-time updates when ballots are collected, received, and accepted by election offices [10]. Similarly, BallotScout integrates with USPS to provide voters with detailed information on their ballots’ whereabouts throughout the election process [7].
There are several limitations of our study that provide insight into topics for future research. First, this analysis considers malicious attacks that occur on a single-day and is limited by a lack of historical data regarding malicious attacks. Future research could diversify attack scenarios and incorporate a broader range of data to enhance the model’s sensitivity to resource constraints and varying electoral processes. Second, future research could account for resource constraints and system congestion. Lastly, we did not examine the likelihood of attack scenarios altering election outcomes. Successful attacks on voting systems compromise the integrity of elections and undermine public trust in political systems, regardless of whether they change the outcomes. Although not all attacks aim to change election results, this remains a significant concern that future research could investigate.